OpenStudy (anonymous):

Use the formula d=rt o find the value of the missing variable d=74 feet, t=37 seconds r= ___ ft/sec

OpenStudy (anonymous):

rearrange to r=d/t r=74 ft /37s

OpenStudy (anonymous):

Id D=RT then.... R=D/T

OpenStudy (anonymous):

ty